# Robust Learning of Multi-index Models via Iterative Subspace Approximation

### v1

@@ 4 Algorithmic Applications of General MIM Learner
In this section, we present applications of our main result. First, in Section 4.1 we present an application of our algorithmic approach to the class of intersections of halfspaces, establishing Theorem 1.9. Second, in Section 4.2, we show that a slight modification of our algorithm can achieve improved performance in the RCN setting, establishing Theorem 1.8.
Additionally, note that our result on learning Multiclass Linear Classifiers, as presented in Theorem 2.1, can also be derived as a corollary of Theorem 3.3, using similar arguments.
@@ 4.1 Learning Intersections of Halfspaces with Adversarial Label Noise
Fix the class $\mathcal{F}_{K,d}$ of intersections of $K$ halfspaces in $\mathbb{R}^{d}$. Let $D$ be a distribution over $\mathbb{R}^{d}\times\{0,1\}$ whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$ and let $\epsilon,\delta\in(0,1),K,d\in\mathbb{Z}_{+}$. Then, Algorithm 3 draws $N=d^{2}2^{\mathrm{poly}(K/\epsilon)}\log(1/\delta)$ i.i.d. samples from $D$, runs in time $\mathrm{poly}(d,N)$, and returns a hypothesis $h$ such that, with probability at least $1-\delta$, it holds that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h(\mathbf{x})\neq y]\leq K\,O(\mathrm{OPT}\log(1/\mathrm{OPT}))+\epsilon$, where $\mathrm{OPT}=\inf_{f\in\mathcal{F}_{K,d}}\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]$.
For this class of Multi-Index Models, we adapt the general algorithm Algorithm 4 to perform polynomial regression with polynomials of degree at most $2$ in Line 4 of Algorithm 4. This degree suffices for learning intersections of halfspaces, as demonstrated in Proposition 4.2. This task can be efficiently reduced to covariance estimation with respect to the Frobenius norm, as the entries of the covariance matrix correspond to the Hermite coefficients. Therefore, we have that it can be performed in polynomial time using $O(d^{2})$ samples. Consequently, the final sample complexity of the algorithm scales with $d^{2}$ rather than $d^{O(1)}$.
The following proposition shows that the class of intersections of halfspaces satisfies Definition 3.2. To prove this result, we leverage that if a function is far from any function defined over a given subspace, then it must also be far from any constant function on a non-trivial fraction of the cylinders restricted to that subspace.
@@
Let $d,K\in\mathbb{Z}_{+},\zeta,\epsilon\in(0,1)$. $\mathcal{F}_{K,d}$ is a class of $(2,\zeta,\epsilon,K,KO(\zeta\log(1/\zeta))+O(\epsilon),\mathrm{poly}(\epsilon/K),O(\sqrt{\log K}))$-well-behaved Multi-Index Models according to Definition 3.2.
@@
Let $D$ be a distribution over $\mathbb{R}^{d}\times\{0,1\}$ such that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]\leq\zeta$. Let $V$ be a subspace of $\mathbb{R}^{d}$, and let $\tau(x)=x\log(1/x)$. Assume that for all functions $g:V\to\{0,1\}$ it holds that $\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq g(\mathbf{x}^{V})]\geq CK\tau(\zeta)+2\epsilon$, for a sufficiently large constant $C$. Hence, the same holds true for the function $h$ defined as $h(\mathbf{x})=\operatorname*{argmin}_{i\in\{\pm 1\}}\mathbf{Pr}_{(\mathbf{x},y)\sim D}[y\neq i\mid\mathbf{x}^{V}]$, since by the definition $h(\mathbf{x})=h(\mathbf{x}^{V})$. Thus, $\operatorname*{\mathbf{Pr}}[h(\mathbf{x})\neq y]\geq(C-1)K\tau(\zeta)+2\epsilon$.
We first show that if the current subspace $V$ is far from optimal, then by splitting the space into cylinders with respect to each point of $V$, the target function is far from being constant on a nontrivial fraction of these cylinders. Specifically, we show the following
@@
With probability at least $\epsilon$ over $\mathbf{x}_{0}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, independent of $\mathbf{x}$, it holds that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq i\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}]>(C-2)K\tau\left(\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}]\right)+\epsilon$ for all $i\in\{\pm 1\}$ and $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}]\leq 1/(C-2)$.
@@
For the sake of contradiction, assume that the set $S$ of $\mathbf{x}_{0}\in\mathbb{R}^{d}$ such that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[y\neq i\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}]>(C-1)K\tau\left(\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}]\right)+\epsilon$ for all $i\in\{\pm 1\}$ has probability mass less than $\epsilon$. Then, we have that
$\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h(\mathbf{x})\neq y]$ $=\operatorname*{\mathbf{E}}_{\mathbf{x}_{0}^{V}}\left[\min_{\hat{y}\in\{\pm 1\}}\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}\left[\hat{y}\neq y\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}\right]\right]$ (3)
$\leq(C-1)K\operatorname*{\mathbf{E}}_{\mathbf{x}_{0}^{V}}\left[\tau\left(\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}\left[f(\mathbf{x})\neq y\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}\right]\right)\right]+2\epsilon$
$\leq(C-1)K\tau(\zeta)+2\epsilon\;,$
where in the last inequality we used Jensen’s inequality since $\tau$ is concave. This leads to a contradiction, and therefore we have that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\mathbf{x}\in S]>\epsilon$. Observe that for each such $\mathbf{x}_{0}$ as $(C-1)K\tau(\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}])+\epsilon\leq\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[y\neq i\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}]\leq 1$, we have that, for a sufficiently large choice of $C$, the probability $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}]$ is less than the constant specified in the statement for all $\mathbf{x}_{0}\in S$. Finally, noting that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[y\neq i\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}]\leq\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq i\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}]+\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}]$ and $\tau(x)\geq x$ for any $x\in(0,1)$ less than a constant gives us the statement. ∎
@@
Fix a point $\mathbf{x}_{0}\in\mathbb{R}^{d}$ that satisfies the statement of Claim 4.3. Let $D_{V^{\perp}}$ be a distribution supported on $V^{\perp}\times\{\pm 1\}$ defined as $D_{V^{\perp}}(\mathbf{x}^{\perp V},y)=D(\mathbf{x},y\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V})$. By the properties of the Gaussian distribution, we have that the $\mathbf{x}$-marginal of $D_{V^{\perp}}$ is a standard Gaussian. Moreover, by the definition of $D_{V^{\perp}}$, we have that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[f^{\prime}(\mathbf{x})\neq i]>(C-2)K\tau(\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[f^{\prime}(\mathbf{x})\neq y])+\epsilon$ for all $i\in\{\pm 1\}$ and $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f^{\prime}(\mathbf{x})\neq y]\leq 1/(C-2)$, where $f^{\prime}(\mathbf{x})=f(\mathbf{x}^{\perp V}+\mathbf{x}_{0}^{V})$. Finally, applying Proposition 4.4 for every intersection of halfspaces $f^{\prime}(\mathbf{x})=f(\mathbf{x}^{\perp V}+\mathbf{x}_{0}^{V})$ for $\mathbf{x}_{0}^{V}$ that satisfies the statement of Claim 4.3 completes the proof of Proposition 4.2. ∎
@@
Below we state and prove a structural result concerning the intersections of halfspaces. This result asserts that if an intersection of halfspaces is far from being constant, then either the first or the second moment exhibits a non-trivial correlation with its weight vectors. This result is crucial for the execution of our algorithm, as it enables us to improve at every iteration.
@@
Let $d\in\mathbb{Z}_{+}$, $\epsilon,\mathrm{OPT}\in(0,1)$, and let $D$ be a distribution over $\mathbb{R}^{d}\times\{0,1\}$ whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$. Let $f$ be a classifier in $\mathcal{F}_{K,d}$ that satisfies $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]=\mathrm{OPT}$, and assume that $f$ can be expressed as $f(\mathbf{x})=\prod_{i\in[K]}\mathds{1}(\mathbf{w}^{(i)}\cdot\mathbf{x}+t_{i}\geq 0)$ with $\mathbf{w}^{(1)},\dots,\mathbf{w}^{(K)}\in\mathbb{R}^{d},t_{1},\dots,t_{K}\in\mathbb{R}$. If it holds that $\min(\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[y\neq 0],\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[y\neq 1])\geq CK\mathrm{OPT}\log(1/\mathrm{OPT})+\epsilon$ and $\mathrm{OPT}\leq 1/C$ for some sufficiently large absolute constant $C$, then there exists some $i\in[K]$ such that
@@
Let $W=\mathrm{span}(\mathbf{w}^{(1)},\dots,\mathbf{w}^{(K)})$ and without loss of generality assume that $\left\|\mathbf{w}^{(i)}\right\|=1$ for all $i\in[K]$, since we can normalize each halfspace without changing the output of $f$. Assume that $\lvert\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[\mathbf{w}^{(i)}\cdot\mathbf{x}y]\rvert\leq p(\epsilon/K)$, for all $i\in[K]$, where $p$ is a sufficiently small universal polynomial (to be quantified later). We show that there exists $i\in[K]$ such that $\lvert\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[((\mathbf{w}^{(i)}\cdot\mathbf{x})^{2}-1)y]\rvert\geq\mathrm{poly}(\epsilon/K)$ for a sufficiently small universal polynomial. We can decompose the second-moment matrix as follows:
$\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[(\mathbf{x}\mathbf{x}^{\top}-\mathbf{I})y]$ $=\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[(\mathbf{x}\mathbf{x}^{\top}-\mathbf{I})f(\mathbf{x})]+\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[(\mathbf{x}\mathbf{x}^{\top}-\mathbf{I})(\mathds{1}(f(\mathbf{x})\neq y=1)-\mathds{1}(y\neq f(\mathbf{x})=1))]$
@@
where $S(\mathbf{x},y)\coloneqq\mathds{1}(y\neq f(\mathbf{x})=1),T(\mathbf{x},y)\coloneqq\mathds{1}(f(\mathbf{x})\neq y=1)$. Note that ${\mathbf{I}_{1}}$ is the second moment of $f(\mathbf{x})$, i.e., the noiseless classifier, and ${\mathbf{I}_{2}}$ is the contribution corresponding to the noise.
@@
Let $\epsilon\in(0,1)$ be less than a sufficiently small universal constant, and let $R:\mathbb{R}^{d}\to[0,1]$ be a function such that $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[R(\mathbf{x})]=\epsilon$. For all unit vectors $\mathbf{v}\in\mathbb{R}^{d}$ the following holds:
@@
    $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[(\mathbf{v}\cdot\mathbf{x})^{2}R(\mathbf{x})]\lesssim\epsilon\log(1/\epsilon)$.
@@
    $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\mathbf{v}\cdot\mathbf{x}R(\mathbf{x})]\lesssim\epsilon\sqrt{\log(1/\epsilon)}$.

### v2

@@ 4 Algorithmic Applications of General MIM Learner
In this section, we highlight selected applications of our main result. First, in Section 4.1 we present an application of our algorithmic approach to the class of intersections of halfspaces, establishing Theorem 1.13. Second, in Section 4.2, we show that a slight modification of our general algorithm can achieve improved complexity in the RCN setting, establishing Theorem 1.11.
Additionally, we note that our result on learning Multiclass Linear Classifiers, as presented in Theorem 2.1, can also be derived as a corollary of Theorem 3.3, using similar arguments.
@@ 4.1 Learning Intersections of Halfspaces with Adversarial Label Noise
Fix the class $\mathcal{F}_{K,d}$ of intersections of $K$ halfspaces in $\mathbb{R}^{d}$. Let $D$ be a distribution over $\mathbb{R}^{d}\times\{0,1\}$ whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$ and let $\epsilon,\delta\in(0,1),K,d\in\mathbb{Z}_{+}$. Then, Algorithm 3 draws $N=d^{2}2^{\mathrm{poly}(K/\epsilon)}\log(1/\delta)$ i.i.d. samples from $D$, runs in time $\mathrm{poly}(N)$, and returns a hypothesis $h$ such that, with probability at least $1-\delta$, it holds that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h(\mathbf{x})\neq y]\leq K\,O(\mathrm{OPT}\log(1/\mathrm{OPT}))+\epsilon$, where $\mathrm{OPT}=\inf_{f\in\mathcal{F}_{K,d}}\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]$.
Proof.
First, note that without loss of generality we can assume that $\mathrm{OPT}\geq\epsilon^{2}$ by randomly flipping an $\epsilon^{2}$ fraction of the labels with probability $1/2$. Therefore, by the mean value theorem we have that
$(\mathrm{OPT}+\epsilon^{2})\log(1/(\mathrm{OPT}+\epsilon^{2})+1)$ $\leq(\mathrm{OPT})\log(1/(\mathrm{OPT})+1)+\epsilon^{2}\log(1/\mathrm{OPT}+1)$
$\leq(\mathrm{OPT})\log(1/(\mathrm{OPT})+1)+\epsilon\;.$
Hence, by applying Proposition 4.2 we have that the class of intersections of $K$-halfspaces is a class of $(2,\mathrm{OPT}+\epsilon,\epsilon,K,KO(\mathrm{OPT}\log(1/\mathrm{OPT}+1)+\epsilon),\mathrm{poly}(\epsilon/K),O(\sqrt{\log K}))$-well-behaved Multi-Index Models.
Thus, by Theorem 3.3 we have that Algorithm 3 draws $N=d^{2}2^{\mathrm{poly}(K/\epsilon)}\log(1/\delta)$ and in $\mathrm{poly}(N)$ time returns a hypothesis $h$ with error $KO(\mathrm{OPT}\log(1/\mathrm{OPT}+1))+O(\epsilon)$. Finally, note that without loss of generality $\mathrm{OPT}\leq 1/2$, since one of the constant classifiers that predict 00 and $1$ can achieve error at most $1/2$. Consequently, we have that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h(\mathbf{x})\neq y]\leq K\,O(\mathrm{OPT}\log(1/\mathrm{OPT}))+O(\epsilon)$, since $x\log(1/x+1)=\Theta(x\log(1/x))$ for all $x\in(0,1/2]$. ∎
For the class of intersections of $K$ halfspaces, we adapt the general algorithm Algorithm 4 to perform polynomial regression with polynomials of degree at most $2$ in Line 4 of Algorithm 4. This degree suffices for learning intersections of halfspaces, as demonstrated in Proposition 4.2. Moreover, the regression task can be efficiently reduced to covariance estimation with respect to the Frobenius norm because the entries of the covariance matrix correspond directly to the Hermite coefficients. Consequently, the regression can be executed in polynomial time using $O(d^{2})$ samples, so that the overall sample complexity of the algorithm scales as $d^{2}$ rather than $d^{O(1)}$.
The following proposition shows that the class of intersections of halfspaces satisfies the regularity conditions in Definition 3.2.
@@
Let $d,K\in\mathbb{Z}_{+},\zeta,\epsilon\in(0,1)$. $\mathcal{F}_{K,d}$ is a class of $(2,\zeta,\epsilon,K,KO(\zeta\log(1/\zeta+1)+\epsilon),\mathrm{poly}(\epsilon/K),O(\sqrt{\log K}))$-well-behaved Multi-Index Models according to Definition 3.2.
@@
Let $D$ be a distribution over $\mathbb{R}^{d}\times\{0,1\}$ ,where the $\mathbf{x}$-marginals of $D$ are standard normal, such that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]\leq\zeta$. Let $V$ be a subspace of $\mathbb{R}^{d}$, and define $\tau(x)=x\log(1/x+1)$. Assume that for every function $g:V\to\{0,1\}$ it holds that $\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq g(\mathbf{x}^{V})]>CK\tau(\zeta)+2\epsilon$, for a sufficiently large constant $C>0$. In particular, consider the function $h(\mathbf{x})=\operatorname*{argmin}_{i\in\{0,1\}}\mathbf{Pr}_{(\mathbf{x},y)\sim D}[y\neq i\mid\mathbf{x}^{V}]$ which depends only on $\mathbf{x}^{V}$ (i.e., $h(\mathbf{x})=h(\mathbf{x}^{V})$). Then the same lower bound applies to $h$, hence $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h(\mathbf{x})\neq y]>(C-4)K\tau(\zeta)+2\epsilon$, since $\tau(x)\geq x/4$ for all $x\in(0,1].$
We first show that if the subspace $V$ is far from optimal, then when we partition the space into cylinders determined by points in $V$, the target function is far from constant on a nontrivial fraction of these cylinders. Specifically, we show the following
@@
With probability at least $\epsilon$ over a random draw of $\mathbf{x}_{0}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ (independent of $\mathbf{x}$), the following holds: for every $i\in\{0,1\}$,
$\mathbf{Pr}_{(\mathbf{x},y)\sim D}[y\neq i\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}]>(C-4)K\,\tau\left(\mathbf{Pr}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}]\right)+\epsilon\;,$
and $\mathbf{Pr}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}]\leq 4/(C-4)$.
@@
Assume for contradiction that the set $U$ that contains $\mathbf{x}_{0}\in\mathbb{R}^{d}$ such that for all $i\in\{0,1\}$ it holds $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[y\neq i\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}]>(C-4)K\tau\left(\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}]\right)+\epsilon$ has probability mass less than $\epsilon$, i.e., $\operatorname*{\mathbf{Pr}}[x\in U]<\epsilon$. Then, by the definition of the piecewise constant classifier we can write
$\mathbf{Pr}_{(\mathbf{x},y)\sim D}[h(\mathbf{x})\neq y]=\operatorname*{\mathbf{E}}_{\mathbf{x}_{0}^{V}}\left[\min_{i\in\{0,1\}}\mathbf{Pr}_{(\mathbf{x},y)\sim D}\left[i\neq y\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}\right]\right].$
Note that for every $\mathbf{x}_{0}^{V}$ not in $U$, it holds that
$\min_{i\in\{0,1\}}\mathbf{Pr}_{(\mathbf{x},y)\sim D}\bigl{[}i\neq y\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}\bigr{]}\leq(C-4)K\,\tau\Bigl{(}\mathbf{Pr}_{(\mathbf{x},y)\sim D}\bigl{[}f(\mathbf{x})\neq y\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}\bigr{]}\Bigr{)}+\epsilon.$
Therefore, by taking the expectation over $\mathbf{x}_{0}^{V}$ and applying Jensen’s inequality (noting that $\tau(x)=x\log(1/x+1)$ is concave for all $x>0$), we obtain
$\mathbf{Pr}_{(\mathbf{x},y)\sim D}[h(\mathbf{x})\neq y]$ $\leq(C-4)K\,\operatorname*{\mathbf{E}}_{\mathbf{x}_{0}^{V}}\Bigl{[}\tau\left(\mathbf{Pr}_{(\mathbf{x},y)\sim D}\left[f(\mathbf{x})\neq y\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}\bigr{]}\right)\right]+\epsilon+\operatorname*{\mathbf{Pr}}[\mathbf{x}^{V}\in U]$
$\leq(C-4)K\,\tau\left(\operatorname*{\mathbf{E}}_{\mathbf{x}_{0}^{V}}\left[\mathbf{Pr}_{(\mathbf{x},y)\sim D}\left[f(\mathbf{x})\neq y\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}\right]\right]\right)+2\epsilon$
$=(C-4)K\,\tau\left(\mathbf{Pr}_{(\mathbf{x},y)\sim D}\left[f(\mathbf{x})\neq y\right]\right)+2\epsilon\;,$
which contradicts the assumed error lower bound for $h$. Therefore, it holds that $\mathbf{Pr}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\mathbf{x}\in U]\geq\epsilon$. Finally, note that for every $\mathbf{x}_{0}$ in $U$ we have, by the definition of $U$,
$(C-4)K\,\tau\left(\mathbf{Pr}_{(\mathbf{x},y)\sim D}\left[f(\mathbf{x})\neq y\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}\right]\right)+\epsilon\leq\mathbf{Pr}_{(\mathbf{x},y)\sim D}\left[y\neq i\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}\right]\leq 1.$
By using that $\tau(x)\geq x/4$ for all $x\in(0,1]$ we deduce the desired claim. ∎
@@
Fix a point $\mathbf{x}_{0}\in\mathbb{R}^{d}$ that satisfies the statement of Claim 4.3. Let $D_{V^{\perp}}$ be a distribution supported on $V^{\perp}\times\{0,1\}$ defined as $D_{V^{\perp}}(\mathbf{x}^{\perp V},y)=D(\mathbf{x},y\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V})$. By the properties of the Gaussian distribution, we have that the $\mathbf{x}$-marginal of $D_{V^{\perp}}$ is a standard Gaussian distribution. Moreover, by the definition of $D_{V^{\perp}}$, we have that for $f^{\prime}(\mathbf{x})=f(\mathbf{x}^{\perp V}+\mathbf{x}_{0}^{V})$ and all $i\in\{0,1\}$ it holds that
$\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[y\neq i]$ $>(C-4)K\tau(\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[f^{\prime}(\mathbf{x})\neq y])+\epsilon$
$\geq((C-4)K/4)\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[f^{\prime}(\mathbf{x})\neq y]\log(1/\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[f^{\prime}(\mathbf{x})\neq y])+\epsilon$
and $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f^{\prime}(\mathbf{x})\neq y]\leq 4/(C-4)$. Finally, applying Proposition 4.4 for every intersection of halfspaces $f^{\prime}(\mathbf{x})=f(\mathbf{x}^{\perp V}+\mathbf{x}_{0}^{V})$ for $\mathbf{x}_{0}^{V}$ that satisfies the statement of Claim 4.3 completes the proof of Proposition 4.2. ∎
@@
Below, we state and prove a structural result for intersections of halfspaces. In particular, we show that if an intersection of halfspaces is far from being a constant function, then either its first moment or its second moment exhibits a non-trivial correlation with the weight vectors that define it. This structural property is crucial for our algorithm, as it guarantees that at every iteration we can identify a direction along which to improve our approximation.
@@
Let $d\in\mathbb{Z}_{+}$, $\epsilon,\mathrm{OPT}\in(0,1)$, and let $D$ be a distribution over $\mathbb{R}^{d}\times\{0,1\}$ whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$. Let $f\in\mathcal{F}_{K,d}$ be a classifier with error $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]=\mathrm{OPT}$, and assume that $f$ can be expressed in the form $f(\mathbf{x})=\prod_{i\in[K]}\mathds{1}(\mathbf{w}^{(i)}\cdot\mathbf{x}+t_{i}\geq 0)$ with $\mathbf{w}^{(1)},\dots,\mathbf{w}^{(K)}\in\mathbb{R}^{d},t_{1},\dots,t_{K}\in\mathbb{R}$. Let $C>0$ be a sufficiently large universal constant. If it holds that $\min(\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[y\neq 0],\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[y\neq 1])\geq CK\mathrm{OPT}\log(1/\mathrm{OPT})+\epsilon$ and $\mathrm{OPT}\leq 1/C$, then there exists some $i\in[K]$ such that
@@
Let $W=\mathrm{span}(\mathbf{w}^{(1)},\dots,\mathbf{w}^{(K)})$ and without loss of generality assume that $\left\|\mathbf{w}^{(i)}\right\|=1$ for all $i\in[K]$, since normalizing the weight vectors does not change the output of $f$. Assume that for all $i\in[K]$ the first moments are small, i.e., $\lvert\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[\mathbf{w}^{(i)}\cdot\mathbf{x}y]\rvert\leq\epsilon p(\epsilon/K)$, where $p$ is a sufficiently small universal polynomial (to be quantified later). We show that there exists $i\in[K]$ such that the second moment is large, i.e., $\lvert\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[((\mathbf{w}^{(i)}\cdot\mathbf{x})^{2}-1)y]\rvert\geq\mathrm{poly}(\epsilon/K)$ for a sufficiently small universal polynomial. We can decompose the second-moment matrix as follows:
$\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[(\mathbf{x}\mathbf{x}^{\top}-\mathbf{I})y]$ $=\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[(\mathbf{x}\mathbf{x}^{\top}-\mathbf{I})f(\mathbf{x})]+\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[(\mathbf{x}\mathbf{x}^{\top}-\mathbf{I})(\mathds{1}(f(\mathbf{x})\neq y,y=1)-\mathds{1}(y\neq f(\mathbf{x}),f(\mathbf{x})=1))]$
@@
where $S(\mathbf{x},y)\coloneqq\mathds{1}(y\neq f(\mathbf{x}),f(\mathbf{x})=1),T(\mathbf{x},y)\coloneqq\mathds{1}(f(\mathbf{x})\neq y,y=1)$. Note that $\mathbf{I}_{1}$ is the second moment of $f(\mathbf{x})$, i.e., the noiseless classifier, and $\mathbf{I}_{2}$ is the contribution corresponding to the noise.
@@
Let $\epsilon\in(0,1)$ less than a sufficiently small universal constant and let $R:\mathbb{R}^{d}\to[0,1]$ be a function such that $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[R(\mathbf{x})]=\epsilon$. Then, for every unit vector $\mathbf{v}\in\mathbb{R}^{d}$, the following bounds hold:
@@
    $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[R(\mathbf{x})(\mathbf{v}\cdot\mathbf{x})^{2}]\lesssim\epsilon\log(1/\epsilon)$.
@@
    $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[R(\mathbf{x})(\mathbf{v}\cdot\mathbf{x})]\lesssim\epsilon\sqrt{\log(1/\epsilon)}$.